OpenStudy (anonymous):

find the diameter of a cone that has a volume of 83.74 cubic inches and a height of 5 inches. use 3.14 for pie 3 inches 4 inches 8 inches 16 inches

OpenStudy (zpupster):

V = (1/3)pi*r^2h diameter = 2r 3V = pi*r*^2h r^2 = 3V/pi*h r = sqrt(3V/pi*h) diameter = 2sqrt(3V/pi*h)